OpenStudy (anonymous):

Match each system of linear equations with the quadrant in which the solution lies.

OpenStudy (anonymous):

7x – 3y = 21 -4x + y = 2 ----------------- x – 2y = -6 4x – 2y = -4 --------------------- -10x – 5y = -10 x + 2y = -14 ------------------- -3x + y = 9 8x + y = -16

OpenStudy (anonymous):

A) I B) II C) IV D) III

OpenStudy (anonymous):

Solve each one of them and see where the solutions are

OpenStudy (anonymous):

This the solution of the first one using wolfram alpha http://www.wolframalpha.com/input/?i=Solve+7x+%E2%80%93+3y+%3D+21+and+-4x+%2B+y+%3D+2

OpenStudy (anonymous):

Both solutions are negative. Which quadrant do they belong too

OpenStudy (anonymous):

1. III 2. I 3. IV 4. II